### v1

@@ 1.1 Condensation Conjectures
Although Langlands conjecture is slightly relatively more representation theoretic in [La], many of the current existing Langlands parametrization happen over certain moduli stacks, such as the group double quotients for the number fields, the group double quotients for the local fields, where the corresponding $\ell$-adic perverse sheaves will capture the action from the motivic Galois groups and the Langlands dual groups, i.e. a map to the Bernstein centers. Condensation is needed in the local setting, and especially when we are using cerain analytic methods, such as Scholze’s diamonds and even the corresponding analytic cohomologies. The current paper is along certain direction generalizing significantly the paper [BS] in the following sense. First we consider finite extension of a local nonarchimedean field $X$, which is assumed to be a finite extension of the $p$-adic number field or a local function field such as $\mathbb{F}_{p}((z))$. Over $X$ there are many considerations can be made beyond the consideration in [BS]. [BS] actually considered certain condensed morphism in the $p$-adic consideration:
@@ 1.2 Generalized Langlands Parameter Stackification
$\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]}$ are $\infty$-stacks over derived rigid analytic spaces $\mathrm{RAS}^{\blacksquare}_{X}$, parametrizing arithmetically all the $(\varphi,\Gamma_{n})$-$G^{\mathrm{Lan}}(\overline{T})$-bundles over the generalized Robba ring $\Pi_{X,\blacksquare}[\log(1+T)^{1/n}]$ in the imperfect setting. We then have the corresponding coherent sheaves over these $\infty$-stacks, where we use the notation $\mathrm{Coh}_{\blacksquare}(\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]})$ to denote the corresponding condesed coherent sheaves.
@@ 1.3 Results
$\mathrm{FilFibCat}_{\mathrm{crys},n}(\square),\mathrm{FilFibCat}_{\mathrm{crys},n,\mathrm{Frob}}(\square),$ (9)
@@ 2 Stackification
$\mathrm{FilFibCat}_{\mathrm{crys},n}(\square),\mathrm{FilFibCat}_{\mathrm{crys},n,\mathrm{Frob}}(\square),$ (17)
@@ 2.1 Stackification for Generalized Hodge Modules
$\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]}$ are $\infty$-stacks over derived rigid analytic spaces $\mathrm{RAS}^{\blacksquare}_{X}$, parametrizing arithmetically all the $(\varphi,\Gamma_{n})$-$G^{\mathrm{Lan}}(\overline{T})$-bundles over the generalized Robba ring $\Pi_{X,\blacksquare}[\log(1+T)^{1/n}]$ in the imperfect setting. We then have the corresponding coherent sheaves over these $\infty$-stacks, where we use the notation $\mathrm{Coh}_{\blacksquare}(\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]})$ to denote the corresponding condesed coherent sheaves.
@@
$\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]}$ are $\infty$-stacks over derived rigid analytic spaces $\mathrm{RAS}^{\blacksquare}_{X}$, parametrizing arithmetically all the $(\varphi,\Gamma_{n})$-$G^{\mathrm{Lan}}(\overline{T})$-bundles over the generalized Robba ring $\Pi_{X,\blacksquare}[\log(1+T)^{1/n}]$ in the imperfect setting. We then have the corresponding coherent sheaves over these $\infty$-stacks, where we use the notation $\mathrm{QuasiCoh}_{\blacksquare}(\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]})$ to denote the corresponding condesed quasicoherent sheaves.
@@ 2.2 Prismatic Consideration
$\mathrm{FilFibCat}_{\mathrm{crys},n}(\square),\mathrm{FilFibCat}_{\mathrm{crys},n,\mathrm{Frob}}(\square).$ (33)
@@
$\mathrm{FilFibCat}_{\mathrm{crys},n}(\square),\mathrm{FilFibCat}_{\mathrm{crys},n,\mathrm{Frob}}(\square),$ (34)
@@ 2.3 $p$-adic Motives
Then by taking the corresponding condesed analytification we have the corresponding analytic version of the stacks.
@@
Then by taking the corresponding condesed analytification we have the corresponding analytic version of the stacks.
@@ 2.4 Construction of Generalized Langlands Correspondence
Then the corresponding push forward along the structure morphism of this stack will generated the desired complex in our setting to realize the corresponding excursion operators as in [FS], which produces the corresponding map from the stack of Langlands paraters cocycle spaces to the Bernstein center.
@@ 3 Generalized geometrized $p$-adic Langlands Program
3.1 Generalized geometrized Breuil-Schneider paramtrization
@@
Then by taking the corresponding condesed analytification we have the corresponding analytic version of the stacks.
@@ 3.2 Generalized Langlands Parameter Stackification
$\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]}$ are $\infty$-stacks over derived rigid analytic spaces $\mathrm{RAS}^{\blacksquare}_{X}$, parametrizing arithmetically all the $(\varphi,\Gamma_{n})$-$G^{\mathrm{Lan}}(\overline{T})$-bundles over the generalized Robba ring $\Pi_{X,\blacksquare}[\log(1+T)^{1/n}]$ in the imperfect setting. We then have the corresponding coherent sheaves over these $\infty$-stacks, where we use the notation $\mathrm{Coh}_{\blacksquare}(\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]})$ to denote the corresponding condesed coherent sheaves.

### v2

@@ 1.1 Condensation Conjectures
Although Langlands conjecture is slightly relatively more representation theoretic in [La], many of the current existing Langlands parametrization happen over certain moduli stacks, such as the group double quotients for the number fields, the group double quotients for the local fields, where the corresponding $\ell$-adic perverse sheaves will capture the action from the motivic Galois groups and the Langlands dual groups, i.e. a map to the Bernstein centers. Condensation is needed in the local setting, and especially when we are using certain analytic methods, such as Scholze’s diamonds and even the corresponding analytic cohomologies. The current paper is along certain direction generalizing significantly the paper [BS] in the following sense. First we consider finite extension of a local nonarchimedean field $X$, which is assumed to be a finite extension of the $p$-adic number field or a local function field such as $\mathbb{F}_{p}((z))$. Over $X$ there are many considerations can be made beyond the consideration in [BS]. [BS] actually considered certain condensed morphism in the $p$-adic consideration:
@@ 1.2 Generalized Langlands Parameter Stackification
$\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]}$ are $\infty$-stacks over derived rigid analytic spaces $\mathrm{RAS}^{\blacksquare}_{X}$, parametrizing arithmetically all the $(\varphi,\Gamma_{n})$-$G^{\mathrm{Lan}}(\overline{T})$-bundles over the generalized Robba ring
$\Pi_{X,\blacksquare}[\log(1+T)^{1/n}]$ (9)
in the imperfect setting. We then have the corresponding coherent sheaves over these $\infty$-stacks, where we use the notation $\mathrm{Coh}_{\blacksquare}(\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]})$ to denote the corresponding condesed coherent sheaves.
@@ 1.3 Results
$\mathrm{FilFibCat}_{\mathrm{crys},n}(\square),$ (10)
$\mathrm{FilFibCat}_{\mathrm{crys},n,\mathrm{Frob}}(\square),$ (11)
@@
Remark 3.
[…]
$\mathrm{Specformal}\Pi_{\mathrm{deRham}}(\sharp\square^{\flat})=\mathrm{Specformal}\mathrm{WittVector}_{\mathcal{O}}(\sharp\square^{\flat})[1/p]_{I_{\sharp\square}}.$ (19)
Then we take the corresponding condensed analytification. If we have a fixed prism as the base ring then we can just use the Cartier divisor for this ring to take the completion with respect to this divisor after inverting $p$. In this manner is $X$ is $z$-adic we can also use the corresponding $\mathrm{WittVector}_{\mathcal{O}}(\square)$ to define the corresponding $z$-adic de Rham period ring which gives rise to $z$-adic prismatization after [BS2], [BLA], [DC]. However here we only use $p$-adic prismatization.
@@ 2 Stackification
$\mathrm{FilFibCat}_{\mathrm{crys},n}(\square),$ (20)
$\mathrm{FilFibCat}_{\mathrm{crys},n,\mathrm{Frob}}(\square),$ (21)
@@ 2.1 Stackification for Generalized Hodge Modules
$\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]}$ are $\infty$-stacks over derived rigid analytic spaces $\mathrm{RAS}^{\blacksquare}_{X}$, parametrizing arithmetically all the $(\varphi,\Gamma_{n})$-$G^{\mathrm{Lan}}(\overline{T})$-bundles over the generalized Robba ring
$\Pi_{X,\blacksquare}[\log(1+T)^{1/n}]$ (29)
in the imperfect setting. We then have the corresponding coherent sheaves over these $\infty$-stacks, where we use the notation $\mathrm{Coh}_{\blacksquare}(\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]})$ to denote the corresponding condesed coherent sheaves.
@@
$\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]}$ are $\infty$-stacks over derived rigid analytic spaces $\mathrm{RAS}^{\blacksquare}_{X}$, parametrizing arithmetically all the $(\varphi,\Gamma_{n})$-$G^{\mathrm{Lan}}(\overline{T})$-bundles over the generalized Robba ring
$\Pi_{X,\blacksquare}[\log(1+T)^{1/n}]$ (30)
in the imperfect setting. We then have the corresponding coherent sheaves over these $\infty$-stacks, where we use the notation $\mathrm{QuasiCoh}_{\blacksquare}(\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]})$ to denote the corresponding condesed quasicoherent sheaves.
@@ 2.2 Prismatic Consideration
$\mathrm{FilFibCat}_{\mathrm{crys},n}(\square),$ (39)
$\mathrm{FilFibCat}_{\mathrm{crys},n,\mathrm{Frob}}(\square).$ (40)
@@
$\mathrm{FilFibCat}_{\mathrm{crys},n}(\square),$ (41)
$\mathrm{FilFibCat}_{\mathrm{crys},n,\mathrm{Frob}}(\square),$ (42)
@@ 2.3 $p$-adic Motives
Then by taking the corresponding condensed analytification we have the corresponding analytic version of the stacks.
@@
Then by taking the corresponding condensed analytification we have the corresponding analytic version of the stacks.
@@
[…]
$\mathrm{Specformal}\Pi_{\mathrm{deRham}}(\sharp\square^{\flat})=\mathrm{Specformal}\mathrm{WittVector}_{\mathcal{O}}(\sharp\square^{\flat})[1/p]_{I_{\sharp\square}}.$ (62)
Then we take the corresponding condensed analytification. Along this way we have many other stacks such as the corresponding crystalline and the corresponding semi-stable ones again after [BS2], [BLA], [DC]:
$\mathrm{Specformal}\Pi_{\mathrm{cristalline}}(\sharp\square^{\flat}),$ (63)
$\mathrm{Specformal}\Pi_{\mathrm{semistable}}(\sharp\square^{\flat}).$ (64)
If we have a fixed prism as the base ring then we can just use the Cartier divisor for this ring to take the completion with respect to this divisor after inverting $p$. In this manner if $X$ is $z$-adic we can also use the corresponding $\mathrm{WittVector}_{\mathcal{O}}(\square)$ to define the corresponding $z$-adic de Rham period ring which gives rise to $z$-adic prismatization after [BS2], [BLA], [DC]. However here we only use $p$-adic prismatization. If $X$ is $z$-adic we can also use the corresponding
$\mathrm{WittVector}_{\mathcal{O}}(\square)$ (65)
to define the corresponding $z$-adic de Rham period ring which gives rise to $z$-adic prismatization after [BS2], [BLA], [DC]. Then we take the corresponding condensed analytification. First we use the formal spectrum of $\mathrm{WittVector}_{\mathcal{O}}(\square)$ to define the $z$-adic prismatization. Then along this way we have many other stacks such as the corresponding crystalline and the corresponding semi-stable ones again after [BS2], [BLA], [DC]:
$\mathrm{Specformal}\Pi_{\mathrm{cristalline}}(\sharp\square^{\flat}),$ (66)
$\mathrm{Specformal}\Pi_{\mathrm{semistable}}(\sharp\square^{\flat}).$ (67)
Then we will then have the corresponding $z$-adic prismatization for any $z$-adic formal scheme. We then use the notation
$\mathrm{Stack}_{\mathrm{prismatic},\mathrm{deRham}},$ (68)
$\mathrm{Stack}_{\mathrm{prismatic},\mathrm{cristalline}},$ (69)
$\mathrm{Stack}_{\mathrm{prismatic},\mathrm{semistable}}$ (70)
to denote the stacks with the condensed-prismatization in both $p$-adic and $z$-adic setting:
$\mathrm{Stack}^{\blacksquare}_{\mathrm{prismatic},\mathrm{deRham}},$ (71)
$\mathrm{Stack}^{\blacksquare}_{\mathrm{prismatic},\mathrm{cristalline}},$ (72)
$\mathrm{Stack}^{\blacksquare}_{\mathrm{prismatic},\mathrm{semistable}}.$ (73)
Then we have the corresponding $\infty$-categories of solid quasicoherent sheaves over these stacks. All above in this definition are defined over $K$, or a formal scheme $S$ over $\mathcal{O}_{X}$, or certainly $z$-adic rigid analytic space $R$ for instance after [SchD] and [SALBRC]. Then we generalize these definitions to the mixed-parity setting after [BS] for instance in de Rham, cristalline and semi-stable situations:
Definition 8.
[…]
$\mathrm{Specformal}\Pi_{\mathrm{cristalline},n}(\sharp\square^{\flat}),$ (74)
$\mathrm{Specformal}\Pi_{\mathrm{semistable},n}(\sharp\square^{\flat}).$ (75)
If we have a fixed prism as the base ring then we can just use the Cartier divisor for this ring to take the completion with respect to this divisor after inverting $p$. In this manner if $X$ is $z$-adic we can also use the corresponding $\mathrm{WittVector}_{\mathcal{O}}(\square)$ to define the corresponding $z$-adic de Rham period ring which gives rise to $z$-adic de Rham prismatization after [BS2], [BLA], [DC]. However here we only use $p$-adic prismatization. If $X$ is $z$-adic we can also use the corresponding $\mathrm{WittVector}_{\mathcal{O}}(\square)$ to define the corresponding $z$-adic de Rham period ring which gives rise to $z$-adic de Rham prismatization after [BS2], [BLA], [DC]. Then we take the corresponding condensed analytification. Then along this way we have the corresponding crystalline and the corresponding semi-stable ones again after [BS2], [BLA], [DC] by adding $n$-th roots of the element $t_{\sharp\square}$:
$\mathrm{Specformal}\Pi_{\mathrm{cristalline},n}(\sharp\square^{\flat}),$ (76)
$\mathrm{Specformal}\Pi_{\mathrm{semistable},n}(\sharp\square^{\flat}).$ (77)
We then use the notation
$\mathrm{Stack}_{\mathrm{prismatic},\mathrm{deRham},n},$ (78)
$\mathrm{Stack}_{\mathrm{prismatic},\mathrm{cristalline},n},$ (79)
$\mathrm{Stack}_{\mathrm{prismatic},\mathrm{semistable},n}$ (80)
to denote the stacks with the condensed-prismatization in both $p$-adic and $z$-adic setting:
$\mathrm{Stack}^{\blacksquare}_{\mathrm{prismatic},\mathrm{deRham},n},$ (81)
$\mathrm{Stack}^{\blacksquare}_{\mathrm{prismatic},\mathrm{cristalline},n},$ (82)
$\mathrm{Stack}^{\blacksquare}_{\mathrm{prismatic},\mathrm{semistable},n}.$ (83)
Then we have the corresponding $\infty$-categories of solid quasicoherent sheaves over these stacks. All above in this definition are defined over $K$.
@@
Proof.
Theorem 16.
Now we assume that the corresponding $X$ is $z$-adic. And we assume the corresponding de Rham stack in this theorem is $z$-adic as well. Over the stack $\mathrm{Stack}_{\mathrm{Bun},G,X,n}$ we have the corresponding $v$-stacks of categories of prismatic quasi-coherent sheaves:
They cover the objects related to lisse $z$-adic representations of the reductive groups [^1] Again we don’t know what is the $z$-adic local Langlands correspondence but we conjecture this is way to define that in a geometrized fashion..
@@ 2.4 Construction of Generalized Langlands Correspondence
Then the corresponding push forward along the structure morphism of this stack will generated the desired complex in our setting to realize the corresponding excursion operators as in [FS], which produces the corresponding map from the stack of Langlands paraters cocycle spaces to the Bernstein center. ∎
@@ 3 Generalized geometrized $p$-adic Langlands Program
3.1 Generalized geometrized Breuil-Schneider parametrization
@@
Then by taking the corresponding condensed analytification we have the corresponding analytic version of the stacks.
@@ 3.2 Generalized Langlands Parameter Stackification
$\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]}$ are $\infty$-stacks over derived rigid analytic spaces $\mathrm{RAS}^{\blacksquare}_{X}$, parametrizing arithmetically all the $(\varphi,\Gamma_{n})$-$G^{\mathrm{Lan}}(\overline{T})$-bundles over the generalized Robba ring
$\Pi_{X,\blacksquare}[\log(1+T)^{1/n}]$ (110)
in the imperfect setting. We then have the corresponding coherent sheaves over these $\infty$-stacks, where we use the notation $\mathrm{Coh}_{\blacksquare}(\mathrm{Stack}^{\blacksquare}_{(\varphi,\Gamma_{n}),G^{\mathrm{Lan}}(\overline{T}),\Pi_{X}[\log(1+T)^{1/n}]})$ to denote the corresponding condesed coherent sheaves.